OpenStudy (anonymous):

please help!!! Add or subtract. Simplify the answer to lowest term (6/r) + (7/(r-4)) is the answer 13/(r2-4)

jhonyy9 (jhonyy9):

(6/r)+(7/(r-4))=(6(r-4)+7r)/(r(r-4))=(6r-24+7r)/(r*2-4r)=(13r-24)/(r*2-4r)

OpenStudy (anonymous):

then = 13/(r2-4)

jhonyy9 (jhonyy9):

(13r/r(r-4))-24/(r(r-4))=(13/(r-4))-(24/(r*2-4r))

jhonyy9 (jhonyy9):

this is the answer

OpenStudy (anonymous):

is totally simplifued??

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

not 13/r^2-4

jhonyy9 (jhonyy9):

is what i have wrote but can tel me how have to you resulted this answer ?

OpenStudy (anonymous):

an equation helper

jhonyy9 (jhonyy9):

so i understand but if you check it this answer not can being correct check please what i have wrote step by step

jhonyy9 (jhonyy9):

and hope so much that you will can understanding all

OpenStudy (anonymous):

thx u